13. A bill that is designed to reduce the nations budget deficit passes both houses of legislature. Congress tells us that the bill will reduce the deficit by $500 billion over 10 years. What it does not tell us is the timing of the reductions as shown below. Deficit Reduction $ 25 Billion $ 30 Billion $ 35 Billion $40 Billion $45 Billion $ 55 Billion $ 60 Billion $ 65 Billion $ 75 Billion $ 85 Billion Year 10 If the federal government can borrow at 8%, what is the true deficit reduction in the bill?
Answer #1

Here we will have to compute the present value of the deficit reduction by discounting the quantum of reduction for each year at a rate of 8% discount rate.

In other words the concept of net present value (NPV) will be used.

Year Deficict reduction 1+r PVIF Present value of deficit reduction
1 25.00 1.08 0.9259 23.15
2 30.00 0.8573 25.72
3 35.00 0.7938 27.78
4 40.00 0.7350 29.40
5 45.00 0.6806 30.63
6 55.00 0.6302 34.66
7 60.00 0.5835 35.01
8 65.00 0.5403 35.12
9 75.00 0.5002 37.52
10 85.00 0.4632 39.37
Total 318.36

Thus we can see that using the concept of time value of money the true deficit reduction in the bill is $318.36 billion.

Calculations:

PVIF = 1/1.08^n where n is the year of cash flow. For instance in year 2 PVIF = 1/1.08^2 = 0.8573

Present value of deficit reduction = actual amount*PVIF. For instance for year 2 present value = 30*0.8573 = 25.72

Total true deficit reduction = sum of all present values as computed above.